General Physics

A trolley while going down an incline plane has an acceleration of 2cm/s . What will be it's velocity 3s after the start

Amisha Rawat
8 Years agoGrade

1 Answer

Arun
8 Years ago

Here, initial velocity, u = 0

 

Acceleration, a = 2 cm s-2

Time, t = 3 s

From, Ist equation of motion, v = u + at = 0 + 2 x 3 = 6 ms-1